OpenStudy (anonymous):

1. solve 3/x+5/x=-4 2.solve 1/3x+9-2/x+3=2 3. solve 8/x+15/x^2=-1

OpenStudy (anonymous):

mutiply through by \(x\) in 1 and 2 and \(x^2\) in 3.

OpenStudy (anonymous):

1. solve 3/x+5/x=-4 x=-2 2.solve 1/3x+9-2/x+3=2 x=-23/6 3. solve 8/x+15/x^2=-1 x=-5,-3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

are u sure they are right

OpenStudy (anonymous):

yes i am:)